A bicycle has a volume of 650 cm3. It is inflated with carbon dioxide to a pressure of 80 psi at 20 C. How many grams of carbon dioxide is contained in the tire?
1
Expert's answer
2021-10-06T13:06:31-0400

Solution.

"V=650cm^3=0.00065m^3;"

"P=80psi=551580Pa;"

"T=293K;"

"M=0.029 kg\/mol;"

"m - ?;"

"PV=\dfrac{m}{M}RT\implies m=\dfrac{PVM}{RT};"


"m=\dfrac{551580\sdot0.00065\sdot0.029}{8.31\sdot293}=0.0043kg=4.3g;"

Answer: "4.3g."